# Spin-up and spin distribution of stellar black holes grown by gas accretion in proto-stellar clusters

### 19 Mar 2026 v1

@@ 3.3 Spin-disk orientation
$\frac{R_{\rm circ}}{r_{\rm g}}=1.2\cdot 10^{4}\left(\tfrac{T_{\rm cl}}{10^{4}{\rm K}}\right)^{-4}\left(\tfrac{m_{\rm BH}}{50\,{\rm M}_{\odot}}\right)^{3}\left(\tfrac{r_{\bullet}}{0.1\,{\rm pc}}\right)^{-3}\left(\tfrac{\omega_{\rm BH}}{\Omega_{{\rm K},\bullet}}\right)^{2}\left(1+\mathcal{M}_{\bullet}^{2}\right)^{-4}$ (30)
@@
where the Mach number is $\mathcal{M}_{\bullet}=v_{\bullet}/c_{{\rm s,cl}}$, the cluster gas temperature is $4\cdot 10^{3}{\rm K}\lesssim T_{\rm cl}(t)\lesssim 1.2\cdot 10^{4}{\rm K}$, and we always assume ionized gas $m_{\rm eff}=0.6m_{\rm p}$ with $\gamma=5/3$. Therefore for smaller BHs, $m_{\rm BH}<10^{3}\,{\rm M}_{\odot}$, the smaller radius is most probably $R_{\rm circ}$, depending however on temperature and BH kinematics (with the exception of strong misalignment when $R_{\rm break}$ is smaller). For larger BHs the $R_{\rm out}$ may be the smallest.
@@
The spin-disk relative orientation depends on $J_{\rm d}/(2J_{\rm BH})$, $\theta_{\rm rel}$, and the BH-alignment timescale $t_{{\rm al},\bullet}$ with respect to the stochastic-kick disruption timestep $\Delta t_{\rm stoch}$ (see Appendix B). As our simulated system evolves, we re-compute $J_{\rm d}(t)$ and $J_{\rm BH}(t)$ dynamically, each time a stochastic kick is applied and then we adopt the following prescription.
@@
  • If $J_{\rm d}/(2J_{\rm BH})<1$, the disk aligns with the BH-spin if the relative inclination satisfies $\theta_{\rm rel}<\tfrac{\pi}{2}$ and counteraligns otherwise (King et al., 2005).
@@
  • If $J_{\rm d}/(2J_{\rm BH})>1$, the BH-spin aligns with the disk for all initial orientations if $t_{{\rm al},\bullet}<\Delta t_{\rm stoch}$. If instead $t_{{\rm al},\bullet}>\Delta t_{\rm stoch}$ we assume steady misalignment (King et al., 2008) and also substitute in eq. (18), $\Lambda=\cos(\theta_{\rm rel})\Lambda_{\rm thin}$ or $\cos(\theta_{\rm rel})\Lambda_{\rm slim}$, depending on disk type.
@@
Once the relative orientation is specified we evolve the spin magnitude (18) for the time interval $\Delta t_{\rm stoch}$ until a new disruption event occurs.
@@
We stress that slim disks, specifically the super-Eddington ones responsible for the high-spin/high-mass BHs, have a significantly shorter BH alignment timescale than thin disks, because the maximum torque is exerted at a shorter distance from the BH. We calculate this quantitatively in detail in Appendix B, with the final expression given in Eq. (40). In practice, the regime $J_{\rm d}/(2J_{\rm BH})>1$, which allows for a net spin-up, may occur when a BH has grown to $m_{\rm BH}>50\,{\rm M}_{\odot}$, accretes at $\dot{m}>1$ (slim disk), and reaches a disk radius of $\mathscr{O}(10^{8}r_{\rm g})$. This follows from equations (24), (29), (30) together with the fact that more massive BHs sink deeper into the center $r_{\bullet}<0.1\,{\rm pc}$. For such BHs we typically get $t_{{\rm al},\bullet}^{\rm slim}\sim 10^{-2}\Delta t_{\rm stoch}$ (see Eq. (40)), and certainly $t_{{\rm al},\bullet}^{\rm slim}\lesssim 0.1\Delta t_{\rm stoch}$ for BHs within the upper BH mass gap range. Therefore, most BHs at and above the mass gap threshold are expected to reach the maximum possible spin allowed by the present accretion geometry, while the more rare cases of $t_{{\rm al},\bullet}>\Delta t_{\rm stoch}$ (and significant $J_{\rm d}$) are already treated through our steady-misalignment prescription. This distinguishes our process from the chaotic accretion for SMBHs (King and Pringle, 2006; King et al., 2008): the underlying disk-formation and evolution physics are different for stellar BHs in clusters and for AGN disks. We note, nevertheless, that strong spin-up (above $a_{*}>0.3$) concerns a minority of the BH population, $\lesssim 10\%$. The majority of the population, corresponding mainly to lower BH masses, stays in the chaotic-accretion regime, and could spin down (those kinematically capable of forming disks) if their initial spin was high. Even so, the transition to high spins at high BH masses and the high-spin/high-mass plateau persists even for higher assumed initial BH spins.
@@
Figure 4: The final BH spin probability density function for $M_{\star}=10^{6}\,{\rm M}_{\odot}$, $r_{c,\star}=1\,{\rm pc}$, star formation efficiency $\varepsilon=0.35$, and ten runs.
@@ 4 Results
We calibrate our primary cluster parameters to be consistent with observations of proto-stellar clusters by JWST in the Cosmic Gems arc galaxy (Adamo et al., 2024). Their observed masses are in the range of $(1-4)\cdot 10^{6}\,{\rm M}_{\odot}$ and the half-light radii are $(0.7-1.7)\,{\rm pc}$. Their observationally estimated age of $\sim 10-50\,{\rm Myr}$ suggests that they possibly are nearly at the final gas expulsion stage, with little or no gas remaining. However, the exact amount of gas remaining is uncertain while the galaxy itself is gas-rich. We assumed a low metallicity of $0.01Z_{\odot}$, in accordance with the low metallicity expected for the Cosmic Gems arc galaxy (Adamo et al., 2024).
@@
We adopted an initial total mass for the initial gas-rich cluster of $M=3\cdot 10^{6}\,{\rm M}_{\odot}$ and a star formation efficiency $\varepsilon=0.35$, which combined give $M_{\star}=10^{6}\,{\rm M}\odot$ for the stellar component, at the low-mass end of the observations. Such a high $\varepsilon$ represents the upper end $\sim 0.3$ of observed star formation efficiencies in embedded clusters (Lada and Lada, 2003; Baumgardt and Kroupa, 2007), appropriate for the extreme densities and low metallicity in such massive, compact proto-stellar clusters as the Cosmic Gems ones. We cover the range of our theoretical Plummer radii of the stellar component at the end of gas depletion equivalent to $r_{c,\star}=1.0-1.3\,{\rm pc}$, that is the appropriate variable to compare with the observed half-light radii. These Plummer radii originate in a more compact iniital gaseous configuration with initial half-mass radii (accounting for both gas and stellar components) $r_{\rm h}^{(\rm ini)}\sim 0.6-0.8\,{\rm pc}$ because of the cluster expansion during gas depletion. This range is ideal since it corresponds to the stall of depletion timescale (Figure 1), with similar $\tau\approx 2.9\,{\rm Myr}$ for the whole range of adopted initial compactness. Such a timescale gives a $99\%$ depletion by $13\,{\rm Myr}$ consistent with the observations that show little to no gas in Cosmic Gems cluster with age $10\,{\rm Myr}$. We terminate simulations at this time of $99\%$ depletion.
@@
Figure 5: Final BH spin with respect to final BH mass for our typical cluster $M_{\star}=10^{6}\,{\rm M}_{\odot}$, $r_{c,\star}=1\,{\rm pc}$, star formation efficiency $\varepsilon=0.35$. (a) Scatter diagram for one indicative run. We have encircled {high-spin, low-mass} and {low-spin, high-mass} outliers. (b) Percentiles, median, and fitted relation, expressed by Eq. (31), for ten runs.
@@
In Figure 3 we depict the scatter diagram of final BH masses with respect to their initial values for one simulation run for visual clarity with $M_{\star}=10^{6}\,{\rm M}_{\odot}$, $r_{c,\star}=1\,{\rm pc}$. We signify the specific BH type based on its formation channel. The PPISN and direct-collapse BHs are those which can populate the upper BH mass gap and reach even IMBH masses, because they are formed earlier originating in more massive stars than the CCSN BHs. We have further investigated in detail the BH mass growth in Paper I, specifying the cluster parameters for which the upper BH mass gap can be populated and IMBH of mass $\sim 10^{3}\,{\rm M}_{\odot}$ can be generated.
@@
A bimodality of spin distribution is evident in Figure 4 with $91\%$ of BHs attaining low spin $a_{*}\leq 0.3$ and $5\%$ high spin $a_{*}\geq 0.7$. We calculate that among all BHs with low spin, the $90\%$ of them have mass $m_{\rm BH}<22.7\,{\rm M}_{\odot}$ and among all BHs with high spin the $90\%$ of them have mass $m_{\rm BH}>63.6\,{\rm M}_{\odot}$ with the maximum median $\bar{a}_{*}$ achieved at $m_{\rm BH}\approx 100\,{\rm M}_{\odot}$.
@@
Figure 5(a) shows the final BH spin corresponding to their final masses for one run. The PPISN BHs of high ZAMS mass progenitor grow in spin more rapidly because they born earlier than all BHs. They present a distinct trend at lower masses $\lesssim 20\,{\rm M}_{\odot}$, however, at higher masses, PPISN and direct-collapse BHs present similar trend with high spins $\gtrsim 0.7$. The CCSN BHs do also spin-up but not above $a_{*}\lesssim 0.7$, although they do not grow significantly in mass. We stress that there exist outliers with low spin at high masses among the Direct-Collapse-high and PPISN-low BHs. On the contrary, the PPISN-high BHs tend to be outliers with high spin and low mass. We discuss outliers in more detail in the followings.
@@
Table 1: Summary of our results for several cluster masses and compactness values. The first column lists the total mass of the stellar component, which originates in a gaseous cluster with total masses $3\cdot 10^{6}$, $10^{6}$, $3\cdot 10^{5}$ $\rm M$$_{\odot}$, respectively, given our adopted star formation efficiency, $\varepsilon=0.35$. Subsequent columns list the stellar component Plummer radius, the initial compactness of the gaseous cluster before gas depletion, the depletion timescale, the maximum BH mass, the maximum BH spin, the fit parameters of the median spin (31), and the last column lists the median saturating spin with its $1\sigma$ percentile for $m_{\rm BH}\geq 100\,{\rm M}_{\odot}$. We chose compactness values that give similar $\tau\lesssim\tau_{\rm stall}$ (Eq. 2) for all cases, while covering the two primary cases of BH masses: either the maximum possible BH mass is generated, $m_{\rm BH}^{\rm max}\approx 10^{3}\,{\rm M}_{\odot}$, or the upper BH mass gap is marginally populated, $m_{\rm BH}^{\rm max}\approx 150\,{\rm M}_{\odot}$. Blue entries indicate stellar cluster mass and size consistent with JWST observations of the Cosmic Gems clusters.
@@
Table 2: Outliers with low spin and high mass. The cluster parameters correspond to the four first rows of Table 1. We denote $\bar{n}$ the number of BHs per cluster within the corresponding mass bin and spin range.
The median $\bar{a}_{*}$ is well fit by a saturating exponential
@@
The parameter $a_{\rm M}$ is the asymptotic median spin value at high masses, and the “transition mass scale” $m_{\rm T}$ designates the point of transition from low to high spin, with $\bar{a}_{*}(m_{\rm T})/a_{\rm M}=0.63$. Remarkably, this function (31) is not specific to pre-assumed cluster mass and size values. It persists for a range of cluster masses and sizes, and not only for intense mass growth with $m_{\rm BH}^{\rm max}\approx 10^{3}\,{\rm M}_{\odot}$, but also even when the upper BH mass gap is only marginally populated with $m_{\rm BH}^{\rm max}\approx 150{\rm M}_{\odot}$. The fit parameter values for several clusters covering those two main cases of BH mass growth is shown in Table 1. The cluster parameters affect the highest possible BH mass and only through this the maximum spin, preserving, however, the general trend of the spin-up equation (31). The saturating median spin at high masses, $m_{\rm BH}\geq 100\,{\rm M}_{\odot}$, ranges from $\bar{a}_{*}^{\rm sat}\sim 0.65$ for low cluster mass, $M_{\star}=10^{5}\,{\rm M}_{\odot}$, to $\bar{a}_{*}^{\rm sat}\sim 0.90$ for high cluster mass, $M_{\star}=10^{6}\,{\rm M}_{\odot}$. The BH transition mass scale ranges slightly about $m_{\rm T}\sim 50\,{\rm M}_{\odot}$ with a median spin at transition $\bar{a}_{*}(m_{\rm T})\sim 0.5$. The fit exponent is $\beta\sim 2.7$, with the exception of the least-massive, least-compact cluster case.
@@
In Figure 5(b) we illustrate the $1\sigma$ and IQR percentiles of the final BH spin with respect to the final mass, together with the median and its fit for the standard parameters $M_{\star}=10^{6}\,{\rm M}_{\odot}$, $r_{c,\star}=1.0\,{\rm pc}$, typical representative values of Cosmic Gems clusters observed by JWST. For heavy BHs, $m_{\rm BH}\geq 100\,{\rm M}_{\odot}$, the $1\sigma$ range spans $a_{*}=0.70-0.96$ with saturating median spin $\bar{a}_{*}^{\rm sat}=0.90$.
@@
These predicted spin values and corresponding masses encompass both components of GW231123 (Abac et al., 2025) with masses $m_{1}=137_{-17}^{+22}\,{\rm M}_{\odot}$, $m_{2}=103_{-52}^{+20}\,{\rm M}_{\odot}$, and spins $a_{*,1}=0.9_{-0.19}^{+0.10}$, $a_{*,2}=0.8_{-0.51}^{+0.20}$. The more massive component has higher spin, consistent with our predicted mass-spin correlation, and both central spin values along with nearly their entire uncertainty ranges lie well within the predicted distribution for their respective masses. This agreement persists down to the less-compact cluster case $r_{c,\star}=1.3\,{\rm pc}$ as in Table 1.
@@
The BHs grown by accretion in our cluster are expected to assemble into merging binary black holes (BBHs) through the standard dynamical channels of dense stellar systems, namely exchange interactions involving existing binaries, three-body binary formation, and subsequent hardening through binary-single and binary-binary encounters (e.g. Benacquista and Downing (2013); Rodriguez et al. (2016)). Later cluster dynamics are expected to randomize spin orientations efficiently. The cosmic evolution of the proto-stellar cluster determines whether the spin-mass correlation of Eq. (31) can survive as an observable imprint on BH spins. If the cluster evolves in the galactic halo into a GC, the imprint is likely to remain largely preserved, because tidal stripping and evaporation (Baumgardt and Makino, 2003) reduce the cluster mass and density, thereby suppressing repeated mergers (Gerosa and Berti, 2019), which could modify the spin-mass relation. By contrast, if the cluster forms in, or later migrates to, the galactic nucleus and becomes part of a nuclear star cluster (NSC), the much higher escape speed can make hierarchical BH mergers efficient (Antonini et al., 2019). The extent to which they wash out the original spin-mass correlation then depends on the cosmic epoch of cluster formation, the delay times for BBH assembly and successive mergers, and the growth history of the host NSC.
@@
We close this section with another result of our analysis. The stochastic nature of accretion allows for outliers with low spin and high mass. In Table 2 we list the number of BHs per cluster of outliers for several mass bins and spin ranges, for the massive clusters of the first four rows of Table 1. For our typical representative of the Cosmic Gems clusters there is expected one mass-gap BH ($\sim 100\,{\rm M}_{\odot}$) per cluster with low spin $a_{*}\sim 0.1$. Massive BHs, $(150-10^{3})\,{\rm M}_{\odot}$, with spin lower than $a_{*}\lesssim 0.3$ are more rare. Clusters with lower mass ($M_{\star}\sim 10^{5}\,{\rm M}_{\odot}$) do not present significant outliers. In addition, as we have noted earlier, the PPISN-high BHs tend to be outliers –sitting outside the $1\sigma$ distribution of Eq. (31)– with higher spin despite their low mass. In particular those outliers with mass $\leq 50{\rm M}_{\odot}$ are well fit by a function, $a_{*,{\rm low-mass}}^{\rm outliers}=A\log{\left(m_{\rm BH}^{\rm fin}\right)}-B,\quad m_{\rm BH}^{\rm fin}\leq 50\,{\rm M}_{\odot}$. The values of those fit parameters and the standard deviation of the spin, $\{A,B,\sigma_{a_{*}}\}$, are for the clusters of the first four rows of Table 1 respectively: $\{0.59,0.35,0.12\}$, $\{0.47,0.28,0.08\}$, $\{0.43,0.17,0.16\}$ and $\{0.39,0.20,0.11\}$.

### 29 Apr 2026 v2

@@ 3.3 Spin-disk orientation
$\frac{R_{\rm circ}}{r_{\rm g}}=1.2\cdot 10^{4}\left(\tfrac{T_{\rm cl}}{10^{4}{\rm K}}\right)^{-4}\left(\tfrac{m_{\rm BH}}{50\,{\rm M}_{\odot}}\right)^{3}\left(\tfrac{r_{\bullet}}{0.1\,{\rm pc}}\right)^{-3}\left(\tfrac{\omega_{\rm BH}}{\Omega_{{\rm K},\bullet}}\right)^{2}\left(1+\mathcal{M}_{\bullet}^{2}\right)^{-4},$ (30)
@@
where the Mach number is $\mathcal{M}_{\bullet}=v_{\bullet}/c_{{\rm s,cl}}$, the cluster gas temperature is $4\cdot 10^{3}{\rm K}\lesssim T_{\rm cl}(t)\lesssim 1.2\cdot 10^{4}{\rm K}$, and we always assume ionized gas, $m_{\rm eff}=0.6m_{\rm p}$, with $\gamma=5/3$. Therefore for smaller BHs, $m_{\rm BH}<10^{3}\,{\rm M}_{\odot}$, the smaller radius is most probably $R_{\rm circ}$, depending, however, on temperature and BH kinematics (with the exception of strong misalignment when $R_{\rm break}$ is smaller). For larger BHs the $R_{\rm out}$ may be the smallest.
@@
The spin-disk relative orientation depends on $J_{\rm d}/(2J_{\rm BH})$, $\theta_{\rm rel}$, and the BH-alignment timescale, $t_{{\rm al},\bullet}$, with respect to the stochastic-kick disruption timestep, $\Delta t_{\rm stoch}$ (see Appendix B). As our simulated system evolved, we recomputed $J_{\rm d}(t)$ and $J_{\rm BH}(t)$ dynamically each time a stochastic kick was applied and then we adopted the following prescription:
@@
  • If $J_{\rm d}/(2J_{\rm BH})<1$, the disk aligns with the BH spin if the relative inclination satisfies $\theta_{\rm rel}<\tfrac{\pi}{2}$ and counteraligns otherwise (King et al., 2005).
@@
  • If $J_{\rm d}/(2J_{\rm BH})>1$, the BH spin aligns with the disk for all initial orientations if $t_{{\rm al},\bullet}<\Delta t_{\rm stoch}$. If instead $t_{{\rm al},\bullet}>\Delta t_{\rm stoch}$ we assume steady misalignment (King et al., 2008) and also substitute in eq. (18), $\Lambda=\cos(\theta_{\rm rel})\Lambda_{\rm thin}$ or $\cos(\theta_{\rm rel})\Lambda_{\rm slim}$, depending on the disk type.
@@
Once the relative orientation was specified, we evolved the spin magnitude (18) for the time interval, $\Delta t_{\rm stoch}$, until a new disruption event occurred.
@@
We stress that slim disks, specifically the super-Eddington ones responsible for the high-spin/high-mass BHs, have a significantly shorter BH alignment timescale than thin disks, because the maximum torque is exerted at a shorter distance from the BH. We calculated this quantitatively in detail in Appendix B, with the final expression given in Eq. (40). In practice, the regime $J_{\rm d}/(2J_{\rm BH})>1$, which allows for a net spin-up, may occur when a BH has grown to $m_{\rm BH}>50\,{\rm M}_{\odot}$, accretes at $\dot{m}>1$ (slim disk), and reaches a disk radius of $\mathscr{O}(10^{8}r_{\rm g})$. This follows from Eqs. (24), (29), and (30), together with the fact that more massive BHs sink deeper into the center, $r_{\bullet}<0.1\,{\rm pc}$. For such BHs we typically get $t_{{\rm al},\bullet}^{\rm slim}\sim 10^{-2}\Delta t_{\rm stoch}$ (see Eq. 40), and certainly $t_{{\rm al},\bullet}^{\rm slim}\lesssim 0.1\Delta t_{\rm stoch}$ for BHs within the upper BH mass gap range. Therefore, most BHs at and above the mass gap threshold are expected to reach the maximum possible spin allowed by the present accretion geometry, while the rarer cases of $t_{{\rm al},\bullet}>\Delta t_{\rm stoch}$ (and significant $J_{\rm d}$) are already treated through our steady-misalignment prescription. This distinguishes our process from the chaotic accretion for SMBHs (King and Pringle, 2006; King et al., 2008): the underlying disk-formation and evolution physics are different for stellar BHs in clusters and for AGN disks. We note, nevertheless, that strong spin-up (above $a_{*}>0.3$) concerns a minority of the BH population, $\lesssim 10\%$. The majority of the population, corresponding mainly to lower BH masses, stays in the chaotic-accretion regime, and could spin down (those kinematically capable of forming disks) if the initial spin was high. Even so, the transition to high spins at high BH masses and the high-spin/high-mass plateau persists even for higher assumed initial BH spins.
@@
Figure 4: Final BH spin probability density function for $M_{\star}=10^{6}\,{\rm M}_{\odot}$, $r_{c,\star}=1\,{\rm pc}$, star formation efficiency $\varepsilon=0.35$, and ten runs.
@@ 4 Results
We calibrated our primary cluster parameters to be consistent with observations of proto-stellar clusters by JWST in the Cosmic Gems arc galaxy (Adamo et al., 2024). Their observed masses are in the range of $(1-4)\cdot 10^{6}\,{\rm M}_{\odot}$ and the half-light radii are $(0.7-1.7)\,{\rm pc}$. Their observationally estimated age of $\sim 10-50\,{\rm Myr}$ suggests that they are possibly nearly at the final gas expulsion stage, with little or no gas remaining. However, the exact amount of gas remaining is uncertain, while the galaxy itself is gas-rich. We assumed a low metallicity of $0.01Z_{\odot}$, in accordance with the low metallicity expected for the Cosmic Gems arc galaxy (Adamo et al., 2024).
@@
We adopted an initial total mass for the initial gas-rich cluster of $M=3\cdot 10^{6}\,{\rm M}_{\odot}$ and a star formation efficiency of $\varepsilon=0.35$, which combined give $M_{\star}=10^{6}\,{\rm M}\odot$ for the stellar component, at the low-mass end of the observations. Such a high $\varepsilon$ represents the upper end, $\sim 0.3$, of observed star formation efficiencies in embedded clusters (Lada and Lada, 2003; Baumgardt and Kroupa, 2007), appropriate for the extreme densities and low metallicities in such massive, compact proto-stellar clusters as the Cosmic Gems ones. We cover the range of our theoretical Plummer radii of the stellar component at the end of gas depletion equivalent to $r_{c,\star}=1.0-1.3\,{\rm pc}$, which is the appropriate variable to compare with the observed half-light radii. These Plummer radii originate in a more compact iniital gaseous configuration with initial half-mass radii (accounting for both gas and stellar components) of $r_{\rm h}^{(\rm ini)}\sim 0.6-0.8\,{\rm pc}$ because of the cluster expansion during gas depletion. This range is ideal, since it corresponds to the stall in the depletion timescale (Fig. 1), with similar $\tau\approx 2.9\,{\rm Myr}$ for the whole range of adopted initial compactness. Such a timescale gives a $99\%$ depletion by $13\,{\rm Myr}$ consistent with the observations that show little to no gas in the Cosmic Gems cluster with an age of $10\,{\rm Myr}$. We terminated the simulations at this time of $99\%$ depletion.
@@
Figure 5: Final BH spin with respect to final BH mass for our typical cluster with $M_{\star}=10^{6}\,{\rm M}_{\odot}$, $r_{c,\star}=1\,{\rm pc}$, and a star formation efficiency of $\varepsilon=0.35$. (a) Scatter diagram for one indicative run. We have circled {high-spin, low-mass} and {low-spin, high-mass} outliers. (b) Percentiles, median, and fit relation, expressed by Eq. (31), for ten runs.
@@
In Fig. 3 we depict the scatter diagram of final BH masses with respect to their initial values for one simulation run for visual clarity with $M_{\star}=10^{6}\,{\rm M}_{\odot}$, $r_{c,\star}=1\,{\rm pc}$. We signify the specific BH type based on its formation channel. The PPISN and direct-collapse BHs are those that can populate the upper BH mass gap and reach even IMBH masses, because they are formed earlier, originating in more massive stars than the CCSN BHs. We have further investigated in detail the BH mass growth in Paper I, specifying the cluster parameters for which the upper BH mass gap can be populated and an IMBH of mass $\sim 10^{3}\,{\rm M}_{\odot}$ can be generated.
@@
A bimodality of spin distribution is evident in Fig. 4, with $91\%$ of BHs attaining a low spin of $a_{*}\leq 0.3$ and $5\%$ a high spin of $a_{*}\geq 0.7$. We calculated that among all of the BHs with a low spin, $90\%$ of them have a mass of $m_{\rm BH}<22.7\,{\rm M}_{\odot}$, and that among all of the BHs with a high spin, $90\%$ of them have a mass of $m_{\rm BH}>63.6\,{\rm M}_{\odot}$, with the maximum median $\bar{a}_{*}$ achieved at $m_{\rm BH}\approx 100\,{\rm M}_{\odot}$.
@@
Figure 5(a) shows the final BH spin corresponding to BH final masses for one run. The PPISN BHs of high ZAMS mass progenitors grow in spin more rapidly because they are born earlier than all BHs. They present a distinct trend at lower masses of $\lesssim 20\,{\rm M}_{\odot}$; however, at higher masses, PPISN and direct-collapse BHs present a similar trend with high spins of $\gtrsim 0.7$. The CCSN BHs do also spin up, but not above $a_{*}\lesssim 0.7$, although they do not grow significantly in mass. We stress that there exist outliers with low spin at high masses among the direct-collapse-high and PPISN-low BHs. On the contrary, the PPISN-high BHs tend to be outliers, with a high spin and a low mass. We discuss outliers in more detail in the following.
Table 1: Fit parameters of BH spin and saturation value for several clusters. [^1] The first column lists the total mass of the stellar component, which originates in a gaseous cluster with total masses of $3\cdot 10^{6}$, $10^{6}$, $3\cdot 10^{5}$ $\rm M$$_{\odot}$, respectively, given our adopted star formation efficiency, $\varepsilon=0.35$. Subsequent columns list the stellar component Plummer radius, the initial compactness of the gaseous cluster before gas depletion, the depletion timescale, the maximum BH mass, the maximum BH spin, and the fit parameters of the median spin (31). The last column lists the median saturating spin with its $1\sigma$ percentile for $m_{\rm BH}\geq 100\,{\rm M}_{\odot}$. We chose compactness values that give similar $\tau\lesssim\tau_{\rm stall}$ (Eq. 2) for all cases, while covering the two primary cases of BH masses: either the maximum possible BH mass, $m_{\rm BH}^{\rm max}\approx 10^{3}\,{\rm M}_{\odot}$, is generated, or the upper BH mass gap is marginally populated, $m_{\rm BH}^{\rm max}\approx 150\,{\rm M}_{\odot}$. Blue entries indicate a stellar cluster mass and size consistent with JWST observations of the Cosmic Gems clusters.
@@
Table 2: Number of high-mass/low-spin BH outliers per cluster for several clusters. [^2] The cluster parameters correspond to the four first rows of Table 1. We denote as $\bar{n}$ the number of BHs per cluster within the corresponding mass bin and spin range.
@@
The median $\bar{a}_{*}$ is well fit by a saturating exponential:
@@
The parameter $a_{\rm M}$ is the asymptotic median spin value at high masses, and the “transition mass scale”, $m_{\rm T}$, designates the point of transition from low to high spin, with $\bar{a}_{*}(m_{\rm T})/a_{\rm M}=0.63$. Remarkably, this function (31) is not specific to pre-assumed cluster mass and size values. It persists for a range of cluster masses and sizes, and not only for intense mass growth with $m_{\rm BH}^{\rm max}\approx 10^{3}\,{\rm M}_{\odot}$, but also even when the upper BH mass gap is only marginally populated, with $m_{\rm BH}^{\rm max}\approx 150{\rm M}_{\odot}$. The fit parameter values for several clusters covering those two main cases of BH mass growth is shown in Table 1. The cluster parameters affect the highest possible BH mass and only through this the maximum spin, preserving, however, the general trend of the spin-up equation, Eq. (31). The saturating median spin at high masses, $m_{\rm BH}\geq 100\,{\rm M}_{\odot}$, ranges from $\bar{a}_{*}^{\rm sat}\sim 0.65$ for a low cluster mass, $M_{\star}=10^{5}\,{\rm M}_{\odot}$, to $\bar{a}_{*}^{\rm sat}\sim 0.90$ for a high cluster mass, $M_{\star}=10^{6}\,{\rm M}_{\odot}$. The BH transition mass scale ranges slightly about $m_{\rm T}\sim 50\,{\rm M}_{\odot}$ with a median spin at transition, $\bar{a}_{*}(m_{\rm T})\sim 0.5$. The fit exponent is $\beta\sim 2.7$, with the exception of the least massive, least compact cluster case.
@@
In Fig. 5(b) we illustrate the $1\sigma$ and IQR percentiles of the final BH spin with respect to the final mass, together with the median and its fit for the standard parameters $M_{\star}=10^{6}\,{\rm M}_{\odot}$ and $r_{c,\star}=1.0\,{\rm pc}$, typical representative values of Cosmic Gems clusters observed by JWST. For heavy BHs, $m_{\rm BH}\geq 100\,{\rm M}_{\odot}$, the $1\sigma$ range spans $a_{*}=0.70-0.96$, with a saturating median spin of $\bar{a}_{*}^{\rm sat}=0.90$.
@@
These predicted spin values and corresponding masses encompass both components of GW231123 (Abac et al., 2025) with masses of $m_{1}=137_{-17}^{+22}\,{\rm M}_{\odot}$ and $m_{2}=103_{-52}^{+20}\,{\rm M}_{\odot}$ and spins of $a_{*,1}=0.9_{-0.19}^{+0.10}$ and $a_{*,2}=0.8_{-0.51}^{+0.20}$. The more massive component has a higher spin, consistent with our predicted mass-spin correlation, and both central spin values along with nearly their entire uncertainty ranges lie well within the predicted distribution for their respective masses. This agreement persists down to the less compact cluster case, $r_{c,\star}=1.3\,{\rm pc}$, as in Table 1.
@@
The BHs grown by accretion in our cluster are expected to assemble into merging binary black holes (BBHs) through the standard dynamical channels of dense stellar systems; namely, exchange interactions involving existing binaries, three-body binary formation, and subsequent hardening through binary-single and binary-binary encounters (e.g., Benacquista and Downing 2013; Rodriguez et al. 2016). Later cluster dynamics are expected to randomize spin orientations efficiently. The cosmic evolution of the proto-stellar cluster determines whether the spin-mass correlation of Eq. (31) can survive as an observable imprint on BH spins. If the cluster evolves in the galactic halo into a GC, the imprint is likely to remain largely preserved, because tidal stripping and evaporation (Baumgardt and Makino, 2003) reduce the cluster mass and density, thereby suppressing repeated mergers (Gerosa and Berti, 2019), which could modify the spin-mass relation. By contrast, if the cluster forms in, or later migrates to, the galactic nucleus and becomes part of a nuclear star cluster (NSC), the much higher escape speed can make hierarchical BH mergers efficient (Antonini et al., 2019). The extent to which they wash out the original spin-mass correlation then depends on the cosmic epoch of cluster formation, the delay times for BBH assembly and successive mergers, and the growth history of the host NSC.
@@
We close this section with another result of our analysis. The stochastic nature of accretion allows for outliers with a low spin and a high mass. In Table 2 we list the number of BHs per cluster of outliers for several mass bins and spin ranges, for the massive clusters of the first four rows of Table 1. For our typical representative of the Cosmic Gems clusters, there is an expected one mass-gap BH ($\sim 100\,{\rm M}_{\odot}$) per cluster with a low spin of $a_{*}\sim 0.1$. Massive BHs, $(150-10^{3})\,{\rm M}_{\odot}$, with a spin lower than $a_{*}\lesssim 0.3$ are rarer. Clusters with a lower mass ($M_{\star}\sim 10^{5}\,{\rm M}_{\odot}$) do not present significant outliers. In addition, as we have noted earlier, the PPISN-high BHs tend to be outliers – sitting outside the $1\sigma$ distribution of Eq. (31) – with higher spin despite their low mass. In particular those outliers with a mass of $\leq 50{\rm M}_{\odot}$ are well fit by a function, $a_{*,{\rm low-mass}}^{\rm outliers}=A\log{\left(m_{\rm BH}^{\rm fin}\right)}-B,\quad m_{\rm BH}^{\rm fin}\leq 50\,{\rm M}_{\odot}$. The values of those fit parameters and the standard deviation of the spin, $\{A,B,\sigma_{a_{*}}\}$, are for the clusters of the first four rows of Table 1, respectively: $\{0.59,0.35,0.12\}$, $\{0.47,0.28,0.08\}$, $\{0.43,0.17,0.16\}$, and $\{0.39,0.20,0.11\}$.